OpenStudy (ilovecake):

The figures shown are similar. What is the scale factor? Help! Medal! A.3 B.1/7 C.2/3 D.1/3

sammixboo (sammixboo):

Scale factor is basically the number you multiply all measurements to get another measurements.

sammixboo (sammixboo):

other*** Anyways, we see that the 21 is changing to 14. What times 21 equals 14? To get that, solve this 21x = 14

OpenStudy (ilovecake):

7?

OpenStudy (ilovecake):

@sammixboo

sammixboo (sammixboo):

Not quite. To solve this, you divde 21 on both sides of the equations 21x/21 = 14/21 x = 14/21 How do you simplify 14/21?

OpenStudy (ilovecake):

2/3

OpenStudy (ilovecake):

So answer is 2/3?

OpenStudy (sissyedgar):

i agree with 2/3

OpenStudy (ilovecake):

So the answer is 2/3?

OpenStudy (ilovecake):

@sammixboo